OpenStudy (anonymous):

PLEEZE HELP WILL FAN AND MEDAL !!!!! Two systems of equations are shown below: System A System B 2x + y = 5 -10x + 19y = -1 -4x + 6y = -2 -4x + 6y = -2 Which of the following statements is correct about the two systems of equations?

OpenStudy (anonymous):

options They will have the same solutions because the first equation of System B is obtained by adding the first equation of System A to 2 times the second equation of System A. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 3 times the second equation of System A. The value of x for System B will be -5 times the value of x for System A because the coefficient of x in the first equation of System B is -5 times the coefficient of x in the first equation of System A. The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding -12 to the first equation of System A and the second equations are identical.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@AlexandervonHumboldt2 @dan815

OpenStudy (mathmale):

I assume this is system A: 2x + y = 5 -10x + 19y = -1 and that this is system B: -4x + 6y = -2 -4x + 6y = -2

OpenStudy (mathmale):

I see immediately that System B presents two identical equations. What does this mean in terms of the math involved?

OpenStudy (mathmale):

Regarding System A: 2x + y = 5 -10x + 19y = -1 Why not actually solve this system as a way to identifying the correct answer? Hint: multiply the first equation (all of it) by 5. Then add this result to the 2nd equation and solve for y.

OpenStudy (anonymous):

System A 2x + y = 5 -4x + 6y = -2 System B -10x + 19y = -1 -4+ 6y = -2

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Sorry, I don't follow what you've shared. Please verify this: System A is as follows: 2x + y = 5 -10x + 19y = -1 If so, mult. the first equation by 5 and solve the resulting system.

OpenStudy (mathmale):

sorry, Layla, I'd like to help, but OpenStudy tells me you're "just looking around."